OpenStudy (anonymous):

An athlete can run long distances at 4 metres per second. How far an she run in an hour???

OpenStudy (anonymous):

3600 * 4 = 14 400 metres = 14,4 km

OpenStudy (anonymous):

hour = 3600 seconds. 4*3600 = 14400 meters

OpenStudy (anonymous):

1 hour=60 minutes 1 minute=60 seconds 1 hour = 60*60=3600 seconds \[distance=speed*time\]

OpenStudy (anonymous):

s = d/t (=) d = s*t = 3600*4

OpenStudy (anonymous):

great guys.